OpenStudy (anonymous):

Given a circle Q where m of arc XY: m of arc YZ: m of arc ZX = 5:6:7 how do I find the measure of angles?

OpenStudy (anonymous):

m of arc XY = 100 degrees m of arc YZ = 120 degrees and m of arc ZX = 140 degrees

OpenStudy (anonymous):

m arc 1 is angle XQZ M angle 2 is angle YQZ m of angle 3 is angle WQZ ??? how do I answer this

OpenStudy (anonymous):

I also know the m of angle 4 is 30 degrees I need to find the measure of angle 5 which is angle XZQ

OpenStudy (anonymous):

Can someone draw this circle and help me? Please

OpenStudy (amistre64):

this one?

OpenStudy (anonymous):

yes please

OpenStudy (amistre64):

itll have to wait; might not be back on. the library is closing up and i have to find another wifi spot...

OpenStudy (anonymous):

ok boo thank you... I'll be here

OpenStudy (amistre64):

how do we modify this?

OpenStudy (anonymous):

thats good... now... ummm let me see

OpenStudy (anonymous):

you have it labeled correctly

OpenStudy (anonymous):

now, draw an internal line in the circle from Y to Z

OpenStudy (anonymous):

an internal line from Z to X. Q is the center point

OpenStudy (anonymous):

A line from Q to Z as well

OpenStudy (anonymous):

A line from Y to Q and a line from X to Q... kind forms a wierd triangle with a radius QZ

OpenStudy (anonymous):

between XY with Q as the center Angle XQY is angle 1, Angle YQZ is angle 2 and Angle ZQX is angle 3

OpenStudy (anonymous):

Angle 4 is the angle Y makes umm... Angle QYZ and angle 5 is XZQ same triangle as Angle 3

OpenStudy (amistre64):

you have a prnt scrn button :)

OpenStudy (anonymous):

Q is the center of the circle...

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

that allows you to take a screen shot of whateves on the monitor; open up a pint program; and paste it in there and modify it :)

OpenStudy (amistre64):

paint program

OpenStudy (anonymous):

ok let me try this.. never did it before

OpenStudy (anonymous):

where is the paint program? I copied and pasted it into word. ??

OpenStudy (amistre64):

start; accessories; paint is usually the path to take

OpenStudy (amistre64):

start; programs; accessories; paint...

OpenStudy (anonymous):

found it one moment

OpenStudy (anonymous):

thats neat

OpenStudy (amistre64):

it is aint it lol

OpenStudy (anonymous):

ok how to I copy and paste it? It looks a little junky but ha first time user... Just as long as you get the point....

OpenStudy (amistre64):

save it to your desktop; then 'attach file' it to the post

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ha mybe I sould have made it bigger

OpenStudy (amistre64):

its good :)

OpenStudy (amistre64):

so whats the question tat goes with this :)

OpenStudy (anonymous):

This is inscribed angles and arcs, I have to find angle 1, 2 , and 3 and 5

OpenStudy (anonymous):

Ha thanks LOL

OpenStudy (anonymous):

Find the measures of those angles

OpenStudy (amistre64):

ang1 = arc.yx

OpenStudy (amistre64):

ang2 = arc.yz

OpenStudy (anonymous):

in the beginning of my problem Arc XY:arc YZ:Arc ZX = 5:6:7

OpenStudy (anonymous):

not sure why proportions are in this

OpenStudy (amistre64):

5 6 7 probably give you a unit measure but yeah, its added detail i think

OpenStudy (amistre64):

ang3 = arc.xz

OpenStudy (anonymous):

yes to angle 1 yes to angle 2

OpenStudy (anonymous):

yep to angle 3

OpenStudy (amistre64):

ang5 = (180-arc.xz)/2

OpenStudy (anonymous):

I know from the back of the book angle 4 is 30 degrees

OpenStudy (anonymous):

it was an odd number :(

OpenStudy (amistre64):

ang4 = (180 - arc.yz)/2 if you wanna check mymethod here

OpenStudy (amistre64):

ang5 = 20?

OpenStudy (amistre64):

180 - 140 = 40/2 = 20 right?

OpenStudy (anonymous):

ummm yes it looks like a smaller angle... I'm trying to see if angle 4 is working out... sorry ... slow motion over here... :)

OpenStudy (amistre64):

180-120 = 60/2 = 30

OpenStudy (anonymous):

where did 140 come from? and how did you know that?

OpenStudy (amistre64):

xy = 100; yz = 120; zx = 140 interior angles at center; 1 = 100; 2 = 120; 3 = 140

OpenStudy (anonymous):

yeah where did those come from... I know 360 is a circle but how did you know to divide those accordingly?

OpenStudy (amistre64):

all triangles from the center are isosoles; if youknow the inside angle the other two are (180-inside)/2 since they are equal angles

OpenStudy (amistre64):

the measure of the arc isnt a length, its a distance rotated; shich is why its in degrees

OpenStudy (anonymous):

k

OpenStudy (amistre64):

an arc length of 90 degrees mens they turned it 90 degrees lol

OpenStudy (amistre64):

once you know the interior central angles; the other angles are equal in size; and all triangles add up to 180 degrees inside them

OpenStudy (amistre64):

180-central angle = whats left over whats left over/2 is the only thing they an be lol

OpenStudy (anonymous):

hey angle 4 comes out right!!!! Oh wow this is complicated stuff!!!

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

what about angle 1, 2 and 3?? is that all in here? LOL

OpenStudy (anonymous):

angle 1 would be 180 - xy/2?

OpenStudy (anonymous):

You are very smart!!!! and thank you for all you have taught me!!!!

OpenStudy (anonymous):

ha the paint too!!!

OpenStudy (amistre64):

ang1 = its arc length that it is in front of:) and youre welcome

OpenStudy (amistre64):

there are no lines to determine angles associated with ang1

OpenStudy (anonymous):

angle 1 comes out 40

OpenStudy (amistre64):

but if you did put a line ther; then yes, that would be how you find them

OpenStudy (amistre64):

yes; the complement angles to 1 would have been 40 :)

OpenStudy (anonymous):

ha oh boy.... its 100

OpenStudy (amistre64):

yes; ang1 is 100 :)

OpenStudy (anonymous):

Ok I should be good I understand, I think thats cause the angle is on the center... not in one of them triangles... right?

OpenStudy (anonymous):

Angle 1I'm talking about...

OpenStudy (anonymous):

Okay... I'll zip the lip!!! LOL thanks!

OpenStudy (amistre64):

right; center angle sweeps out an arc length of # degrees that arc length is

OpenStudy (amistre64):

if arc length = 120 degrees, then angle has to be 120 degrees

OpenStudy (anonymous):

completely makes sense!

OpenStudy (anonymous):

you are a rock star!!!

OpenStudy (amistre64):

:) thnx